Mission Status Report #8 Star Date: July 1, 1999
Photo: Science Team Lead, Bill Oegerle accepts the "key to FUSE" from Jim Moore, NASA/GSFC Director of Flight Projects, at Launch Day ceremony at JHU. (Photo courtesy of NASA/GSFC photographer Debbie McCallum). (Click photo to see larger version.)
FUSE CHECKOUT CONTINUES ON TRACK
FUSE has been in orbit 1 week now, and no major problems have been encountered! Mission Status Reports on the web page will start being more sporadic now, when news-worthy events occur (but at least weekly!) for the time being.
Date: Thu, 1 Jul 1999 20:33:04 -0400 (EDT) From: Bill Oegerle Subject: FUSE IOC, July 1 FUSE IOC news, July 1, 1999 Yesterday's activities concluded with the successful turn-on of the FPAEs (focal plane assembly electronics boxes), and successful conclusion of the bright object sensor - bright earth test. The BOS did not trigger when the boresight passed over the earth. We then slewed to the southern continuous viewing zone (CVZ), where we will spend most of our time over the next few weeks). Almost everything is turned on at his point, and all continues to go very well. Today was spent checking out FES-B. Tests included an ADC reset procedure, a decontamination heater test, a TEC test, and several imaging tests. The thermo-electric cooler (TEC) was set at -40 deg C. Two full field of view images were obtained, binned 1x1 and 2x2. These images showed the expected dark current level, as well as a few cosmic rays. During execution of the image-taking script, a script error was encountered (now understood) which resulted in FES-B needing to be rebooted. This was performed without problem. The image-taking script was re-executed later this afternoon without error. We are experiencing some telemetry drop-outs at the Hawaii ground station, which may (or may not) be related to FUSE's antenna null pattern. This null pattern is due to interference between FUSE's two omni antenna beams. So, some time was spent today analyzing the downlink signal to Hawaii at different satellite orientations.
Reported by: Bill Oegerle, Science Ops Lead
Return to the List of Status Reports